# Circuit-tuning: A Mechanistic Approach for Identifying Parameter Redundancy and Fine-tuning Neural Networks

### v2

@@ 6 Conclusion
We extend the concept of intrinsic dimensionality to the node level, and describe the learning process inside a model from the mechanistic view. Further, we propose circuit-tuning as a promising tool for interpretable fine-tuning. We analyze the mechanism behind learning and provide new insights for the learning process. Limitations and future discussions of our work are shown in Appendix A.
@@ Appendix A Limitations and future discussions
  2. 2.
    We only focus on attention heads as most people do. In fact, we believe that a lot of treasures are hidden inside the MLP layers. Recently Ameisen et al. [2], Lindsey et al. [32] applied transcoders to MLP layers and discovered the circuits inside a model. Thus we look forward to seeing circuit-tuning combined with other model components besides attention.
  3. 3.
    We believe that the experiment on complex tasks could be further studied. For example, we could analyze the learning dynamics during the fine-tuning process for COT-based tasks, or COT-free tasks like the gender de-biasing task. However, due to the complexity of these tasks and limited space, we think it is a better idea to study them in a separate work and present them in a single paper. Thus we leave this part for further research.
  4. 4.
    The infra of our method could be improved. We believe it is possible to apply our method to larger models, while it requires stronger frameworks. We leave it for furture development.
Appendix B Social impact
This paper presents work whose goal is to advance the field of Machine Learning. Since our work concerns AI alignment and involves modifying a model’s gender stereotype in Section 5.2, we expect our work to be used under the guidance of human values.
Appendix C Transformer Architecture
@@
A single input of Transformer is $x_{0}\in\mathbb{R}^{T}$, where $T$ is the length of the sequence. The input is firstly embedded into a vector $x\in\mathbb{R}^{D\times T}$ via the embedding matrix $W_{E}\in\mathbb{R}^{D\times V}$, where $D$ is the model dimension. Then $x$ will go through $l$ layers of Transformer blocks for various processings. From the view of [13], we can think of the residual stream as a communication channel that simply receives the output of the self-attention and MLP operations. Each operation reads information from the residual stream and writes the processed information into it. Thus the residual stream is actually the linear sum of various transformations of $x$ together with the original input $x$.
@@
For the MLP operation in each layer, the input $x$ is projected into $x_{in}^{i}\in\mathbb{R}^{D_{mlp}\times T}$ via $W_{in}\in\mathbb{R}^{D_{mlp}\times D}$, and projected back to $MLP_{i}(x)\in\mathbb{R}^{D\times T}$ via $W_{out}\in\mathbb{R}^{D\times D_{mlp}}$. In the Llama architecture [59], the input $x$ is firstly projected into $x_{pre}^{i}\in\mathbb{R}^{D_{mlp}\times T}$ via $W_{gate}^{i}\in\mathbb{R}^{D_{mlp}\times D}$ and is applied with an activation layer, then a dot product is performed between the activations and $x_{in}^{i}\in\mathbb{R}^{D_{mlp}\times T}$ which is the input projected by $W_{in}\in\mathbb{R}^{D_{mlp}\times D}$. Note that we can also split the MLP into MLP heads, which is done on Llama series models in the complex tasks in our experiments. For details, please refer to Appendix G.2.
@@
Appendix D The Derivation of Edge Contribution
Similar to the definition of attribution score in [56], we can define the contribution of an edge $e:n_{1}\rightarrow n_{2}$ with $n_{1}$ as the upstream node and $n_{2}$ as the downstream node. We use the indirect effect $IE$ to measure the change in the output caused by the patching of the edge. Thus given a dataset $\mathcal{X}$ for patching, the contribution of edge $e$ can be expressed as follows:
@@ Appendix D The Derivation of Edge Contribution
$=\mathbb{E}_{x_{i}\sim\mathcal{X}}\big{[}|IE(e;x_{i})|\big{]}$ (4)
Note that the contribution of edge $e$ is directly reflected in the change of the final output (the logit of the language model, etc), which is the indirect effect caused by the change of the value in the downstream node $n_{2}$, while the change of the node $n_{2}$ is directly caused by the change of the upstream node $n_{1}$. The difference between the direct effect and the indirect effect is that the former keeps all the other nodes that could influence $n_{2}$ unchanged and only studies the influence from $n_{1}$ to $n_{2}$, while the latter allows all the changes in nodes between $n_{2}$ and the logit. For more refined definitions for these two concepts, please refer to [44].
@@
To further simplify Equation D, we apply another Taylor expansion at $n_{1}=n_{1}(x)$ to $n_{2}^{n_{1}}$. Then we have
@@
Thus we come to the final form of edge contribution in Equation D. Our derivation takes the simplest situation into consideration, while it works well in practice. For more discussions on the relevant topic, please refer to [27]. For implementation details, please refer to our code.
Appendix E The proof for the claim in Section 3.4
In Section 3.4, we state that the estimation of the intrinsic dimensionality can be converted to the process of detecting the parameter redundancy in a computational graph. Specifically, we believe that there are mainly two types of redundancies in a model: representation redundancy and forward redundancy. Representation redundancy describes the redundancy of node in representing the features of an input, while forward redundancy describes the redundancy of node in the forward propagation. Thus the redundancy of a node is the combination of the above two types of redundancies, which finally points to the redundancy in parameters.
To prove this, we first propose the definition of representation redundancy and forward redundancy in Appendix E.2 and extend them to the node level in Appendix E.3 and E.4, respectively. Then we prove that the approximation of the indirect effect in Equation (3) consists of two parts that correspond to the representation redundancy and forward redundancy respectively in Appendix E.5 and E.6.
E.1 Preliminaries
Features
Given a representation space $\mathbb{V}^{D}(D=m)$ with a set of bases $\mathbf{E}=\{\mathbf{e_{1},e_{2},...,e_{D}}\}$ where each basis $\mathbf{e_{i}}$ corresponds to a neuron, the feature vector $\mathbf{v_{f}}$ of feature $f$ is a direction in $\mathbb{V}^{D}$:
$\mathbf{v_{f}}=\sum\nolimits_{i=1}^{m}c_{i}\mathbf{e_{i}}=c_{1}\mathbf{e_{1}}+c_{2}\mathbf{e_{2}}+...+c_{m}\mathbf{e_{m}}$ (12)
where the coefficient $c_{i}$ satisfies $|ci|\leq 1$ and $\sum_{i=1}^{m}ci^{2}=1$.
The idea of feature vector follows the linear representation hypothesis that features are represented as vectors in a linear space, the evidence of which has been widely discussed [36, 5, 43].
Superposition
Given a layer in a neural network with its representation space $\mathbb{V}^{D}$, if all features represented in that layer is $\mathcal{F}$, then we have $D<|\mathcal{F}|$, where $|\mathcal{F}|$ is the number of elements in $\mathcal{F}$.
The superposition hypothesis discussed in [41, 12] aims to explain the phenomenon of polysemanticity that one neuron in a neural network often responds to multiple unrelated inputs. Since there are more features than the dimensions, the model learns to compress more features than it has dimensions.
Representation
Considering the settings of feature vector and superposition, the representation $H$ of an input $x$ in $\mathbb{V}^{D}(D=m)$ is the linear combination of its feature vectors:
$H=\sum\nolimits_{i=1}^{t}\mathbf{v_{f_{i}}}=(h_{1},...,h_{m})^{\top},h_{i}=\Big{(}\sum\nolimits_{i=1}^{t}a_{f_{i}}\mathbf{v_{f_{i}}}\Big{)}\cdot\mathbf{e_{i}}$ (13)
where the features of $x$ are $\mathcal{F}_{x}=\{f_{1},f_{2},...,f_{t}\}$ with their magnitudes $A_{x}=\{a_{f_{1}},a_{f_{2}},...,a_{f_{t}}\}$. Thus the activation $h_{i}$ of dimension $i$ in $\mathbb{V}^{D}$ often originates from more than one features.
E.2 Two types of redundancy
Intrinsic dimensionality describes the redundancy in parameters in terms of a specific task, and our goal is to specify it. In this section, we define two types of redundancy from the mechanistic view.
When inspecting a layer with its weight matrix $W^{pre}\in\mathbb{R}^{m\times n}$ that maps an input from $\mathbb{R}^{n}$ to $\mathbb{R}^{m}$, we focus on a representation $H=({h_{1},h_{2},...,h_{m}})^{\top}\in\mathbb{V}^{D}$ in the representation space $\mathbb{V}^{D}(D=m)$ of that layer.
Assumption 1: Feature sparsity
Suppose all features represented in a layer is $\mathcal{F}$. Given a task $T$ with $\mathcal{X}=\{x_{1},x_{2},...,x_{t}\}\sim\mathcal{D}_{T}$ which consists of $t$ samples that follows a specific data distribution $\mathcal{D}_{T}$, the features for representing $\mathcal{X}$ is $\mathcal{F}_{T}\subsetneqq\mathcal{F}$.
This means that when conditioned on a task $T$, the number of features needed is smaller than that of the elements in the universal set $\mathcal{F}$. The assumption follows the concept of feature sparsity [12] that many features do not frequently appear.
Definition 1: Representation redundancy
Given a task $T$ with features $\mathcal{F}_{T}$, we focus on a representation $H=(h_{1},...,h_{m})^{\top}$ and the $i$-th dimension in $\mathbb{V}^{D}(D=m)$. We set the value of the activation $h_{i}$ in that dimension to zero. For each feature $f\in\mathcal{F}_{T}$, we can get the deformed feature vector $\tilde{\mathbf{v_{f}}}$. Given a metric $\Phi$ for measuring the similarity between two vectors and a threshold $\tau_{r}\in(0,1)$ as the tolerance of deformation, if the intervention on the $i$-th dimension satisfies:
$\mathop{min}\limits_{1\leq i\leq|\mathcal{F}_{T}|}\ \big{[}(\Phi(\mathbf{v_{f_{1}}},\tilde{\mathbf{v_{f_{1}}}})),...,\Phi(\mathbf{v_{f_{i}}},\tilde{\mathbf{v_{f_{i}}}}))\big{]}>\tau_{r}$ (14)
then the $i$-th dimension is redundant for representing features $\mathcal{F}_{T}$, and $W_{i,:}^{pre}$ that is the $i$-th column of $W^{pre}$ is a redundant parameter in the sense of representing features.
The idea is that if all features in task $T$ preserve their semantics after perturbation, then the influence from the intervention on the $i$-th dimension to the semantic representation of $\mathcal{F}_{T}$ can be ignored. It is a nice property akin to the robustness of population encoding to noise. Since information is encoded across many cells, a perturbation on a few cells will not destroy the representation [45].
If the representation $H=(h_{1},h_{2},...,h_{m})^{\top}$ is followed by another weight matrix $W^{post}\in\mathbb{R}^{l\times m}$, then the following representation is $H^{\prime}=W^{post}H\in\mathbb{V}^{D^{\prime}}(D^{\prime}=l)$. To study the influence of $H$ on $H^{\prime}$, we split $H^{\prime}$ into the sum of influences from the dimensions in $H$:
$W^{post}H=\sum\nolimits_{i=1}^{m}W_{:,i}^{post}h_{i}$ (15)
Equation (15) shows that the influence from activation $h_{i}$ to $H^{\prime}$ is subject to $W_{:,i}^{post}$. If the magnitude of the weights in $W_{:,i}^{post}$ is small, then the influence from $h_{i}$ will be weakened, leading to a reduced impact on the final result in the follow-up computations. Thus, we provide the following definition:
Definition 2: Forward redundancy
Given a lower bound $\tau_{f}\in(0,1)$, we inspect the redundancy of the $i$-th dimension in $\mathbb{V}^{D}$ in the forward propagation. If the activation $h_{i}$ at that dimension satisfies:
$\bigg{[}\frac{W_{:,i}^{post}h_{i}\cdot H^{\prime}}{||H^{\prime}||}\bigg{]}_{norm}<\tau_{f}$ (16)
then dimension $i$ is redundant, and $W_{i,:}^{pre}$ that is the $i$-th column of $W^{pre}$ is a redundant parameter when considered in the forward propagation.
On the left side of (16) is the normalized projection of $W_{:,i}^{post}h_{i}$ on $H^{\prime}$ among the values in $H$, which measures the weight of activation $h_{i}$ in the following representation $H^{\prime}$, i.e., its influence in the forward propagation.
E.3 The representation redundancy at the node level
The definition of representation redundancy at the neuron level is shown in Section E.2, in which we set the activation value of a dimension to zero and check its redundancy through the semantic change in feature vectors. Specifically, considering an input with its features $\mathcal{F}_{T}=\{f_{1},...,f_{|\mathcal{F}_{T}|}\}$, the determination of the representation redundancy of a node is:
$\mathop{min}\limits_{1\leq i\leq|\mathcal{F}_{T}|}\ \big{[}(\Phi(\mathbf{v_{f_{1}}},\tilde{\mathbf{v_{f_{1}}}})),...,\Phi(\mathbf{v_{f_{i}}},\tilde{\mathbf{v_{f_{i}}}}))\big{]}>\tau_{r}$ (17)
where $\mathbf{v_{f_{i}}}$ and $\tilde{\mathbf{v_{f_{i}}}}$ are the original and deformed feature vectors of the feature $f_{i}$, respectively. The metric $\Phi$ is used to measure the similarity between two vectors (e.g. cosine similarity), and $\tau_{r}\in(0,1)$ is the tolerance for deformation.
In this section, we discuss about the representation redundancy at the node level, i.e., each node consists of a group of neurons, or simply an activation (e.g. the output of an attention head). We define each node $n$ as a group of $N$ neurons, i.e., $n=(h_{1},...,h_{N})^{\top}$. The size of each node is at most an representation in a layer (e.g. a representation in the residual stream). Though a node could also be a layer in a model, we require $N\leq D$ to cover most of the cases for circuit discovery and for convenience of discussions. In this situation, the representation $H$ is a set of nodes $\mathcal{N}=(n_{1},...,n_{|\mathcal{N}|})$, where $|\mathcal{N}|$ is the number of nodes in $H$.
Therefore, the representation redundancy defined at the node level is almost the same as the original definition at the neuron level in section E.2, except that we focus on the redundancy of a node, instead of a single dimension. Hence, when inspecting the representation redundancy of a node $n_{i}$, we set the value of it to a zero vector and check the change in feature vectors. If Equation 17 is satisfied, then the node $n_{i}$ is redundant for representing the features $\mathcal{F}_{T}$, and thus $W_{s_{i},:}^{pre}$ is redundant, where $s_{i}=[(i-1)\cdot N+1,i\cdot N]$ is the slice in $W^{pre}$ for the node $n_{i}\in\mathcal{N}$.
E.4 The forward redundancy at the node level
The definition of forward redundancy at the neuron level is shown in Section E.2. In this section, we discuss about the forward redundancy at the node level.
Consider a representation $H$ in the representation space $\mathbb{V}^{D}(D=m)$ and a follow-up representation $H^{\prime}$ in the representation space $\mathbb{V}^{D^{\prime}}(D^{\prime}=l)$, the relation between $H$ and $H^{\prime}$ is:
$H^{\prime}=W^{post}H=\sum_{i=1}^{m}W_{:,i}^{post}h_{i}$ (18)
where $W^{post}$ is the parameter matrix in the representation space $\mathbb{V}^{D^{\prime}}$. In Equation 18, the representation $H^{\prime}$ is split into the sum of the productions between each neuron $h_{i}$ in $H$ and its corresponding weight $W_{:,i}$, the $i$-th column of $W^{post}$.
Same as the settings in Appendix E.3, we define each node $n$ as a group of $N$ neurons, i.e., $n=(h_{1},...,h_{N})^{\top}$. Thus the representation $H$ is a set of nodes $\mathcal{N}=(n_{1},...,n_{|\mathcal{N}|})$. Similarly, the representation $H^{\prime}$ can be written as $\mathcal{N^{\prime}}=(n_{1}^{\prime},...,n_{|\mathcal{N^{\prime}}|}^{\prime})$.
Therefore, the relation between $H$ and $H^{\prime}$ can be written as:
$H^{\prime}=W^{post}H=\sum_{i=1}^{|\mathcal{N}|}W_{:,s_{i}}^{post}n_{i}$ (19)
where $s_{i}=[(i-1)\cdot N+1,i\cdot N]$ is the slice in $W^{post}$ for the node $n_{i}\in\mathcal{N}$. Thus Equation (16) in the definition of forward redundancy can be extended to the node level:
$\bigg{[}\frac{W_{:,s_{i}}^{post}n_{i}\cdot H^{\prime}}{||H^{\prime}||}\bigg{]}_{norm}<\tau_{f}$ (20)
If Equation 20 is satisfied, then the node $n_{i}$ is redundant in the forward propagation. Thus $W_{s_{i},:}^{pre}$ is redundant when considered in the forward propagation, where $s_{i}=[(i-1)\cdot N+1,i\cdot N]$ is the slice in $W^{pre}$ for the node $n_{i}\in\mathcal{N}$.
E.5 The representation redundancy in the approximation of indirect effect
The first part of the approximation of the indirect effect in Equation (3) is $n(x^{\prime})-n(x)$, where $n(x^{\prime})$ is often the averaged result of all other nodes. Since we consider the absolute value of this term when computing the overall node contribution described in Equation 2, we check the value of $n(x)-n(x^{\prime})$ instead for convenience. The term $n(x)-n(x^{\prime})$ is the difference between the original activation and the averaged activation, and it can be viewed as a centered activation $n(x)^{c}$.
To prove that the first part corresponds to the representation redundancy, we try to confirm that $n(x)^{c}$ is an indicator for representation redundancy. According to the approximation of the indirect effect in Equation 3 which is also the detection for the redundancy in a node, the smaller $n(x)^{c}$ is, the more likely that the node is redundant. Thus we expect that a smaller $n(x)^{c}$ to meet the requirement of representation redundancy.
When $n(x)^{c}$ approaches zero, the activation value $n(x)$ of the node $n$ approaches the mean value $\overline{n(x)}$ of all the nodes. Thus we need to prove that when $n(x)\rightarrow\overline{n(x)}$, the node $n$ is more likely to meet the requirement of the representation redundancy.
According to the assumption of feature sparsity in Section E.2, many features in a representation space do not frequently appear. Specifically, if all the features in the representation space $\mathbb{V}^{D}$ is $\mathcal{F}$, then the features $\mathcal{F}_{T}$ of an input satisfies $\mathcal{F}_{T}\subsetneqq\mathcal{F}$. When the model gets larger with more dimensions, then the sparsity becomes much stronger, then we have $|\mathcal{F}_{T}|\ll|\mathcal{F}|$. The sparsity in features results in a long-tailed distribution of activation values in a representation, as demonstrated in [12]. Therefore, the average value of the activations in a representation could be quite small:
$\overline{n(x)}=\mathcal{O}\left(\frac{1}{D}\right)$ (21)
where $D$ is the number of dimensions in the representation space $\mathbb{V}^{D}$. Hence the activation value (or norm) $n(x)$ of the node $n$ is small when it approaches $\overline{n(x)}$. This implies that few features depend on that node: According to Equation (13), the activation of a dimension is the sum of the projections from all the feature vectors. When it comes to the node level, considering a node $n_{i}=(h_{1},...,h_{N})^{\top}$ which includes $N$ dimensions in the representation, we have:
$h_{i}=\Big{(}\sum\nolimits_{t=1}^{|\mathcal{F}_{T}|}a_{{}_{t}}\mathbf{v_{f_{t}}}\Big{)}\cdot\mathbf{e_{i}},\quad i=1,...,N$ (22)
A small norm of $n_{i}$ implies small activation values $h_{i}(i=1,...,N)$, which means the sum of the projections of the feature vectors is small, and thus the node $n_{i}$ is more likely to be a redundant node for representing the features in $\mathcal{F}_{T}$. To prove this, we introduce $P_{r}$ as the probability that the node $n_{i}$ is redundant. According to the definition of representation redundancy in Appendix E.3, we have:
$P_{r}=\prod_{t=1}^{|\mathcal{F}_{T}|}P[\Phi(\mathbf{v_{f_{t}}},\tilde{\mathbf{v_{f_{t}}}})>\tau_{r}]$ (23)
For each feature $f_{t}\in\mathcal{F}_{T}$, the smaller the projection of its feature vector $\mathbf{v_{f_{t}}}$ on the dimensions in node $n_{i}$, the larger the similarity $\Phi(\mathbf{v_{f_{t}}},\tilde{\mathbf{v_{f_{t}}}})$ between $\mathbf{v_{f_{t}}}$ and $\tilde{\mathbf{v_{f_{t}}}}$, and thus the higher the probability $P[\Phi(\mathbf{v_{f_{t}}},\tilde{\mathbf{v_{f_{t}}}})>\tau_{r}]$ is when all the features in $\mathcal{F}_{T}$ are considered.
Therefore, we can infer that when $n_{i}$ approaches the mean value of all the nodes which is a small value due to feature sparsity, few features rely on the dimensions in $n_{i}$ for representation, and thus the probability $P_{r}$ that $n_{i}$ is redundant is higher. Thus the centered value $n(x)^{c}$ of a node $n$ serves as an indicator for the representation redundancy of that node.
E.6 The forward redundancy in the approximation of indirect effect
The second part of the approximation of the indirect effect in Equation (3) is:
$\nabla_{n}\mathcal{L}_{m}[M(x)]\big{|}_{n(x)}$ (24)
which is the derivative of the metric $\mathcal{L}_{m}$ w.r.t the value of the node $n$.
Now we inspect a node $n_{i}$ in the representation $H$. According to the chain rule and the discussions in Appendix E.4, the derivative in (24) can be written as
$\nabla_{n_{i}}\mathcal{L}_{m}=\sum_{j=1}^{|\mathcal{N^{\prime}}|}\nabla_{n_{j}^{\prime}}\mathcal{L}_{m}\cdot\nabla_{n_{i}}n_{j}^{\prime}$ (25)
where the node $n_{j}^{\prime}(j=1,...,|\mathcal{N}^{\prime}|)$ is the nodes in the follow-up representation, and we have:
$n_{j}^{\prime}=\sum_{i=1}^{|\mathcal{N}|}W_{s_{j},s_{i}}^{post}\cdot n_{i}$ (26)
where $s_{j}=[(j-1)\cdot N+1,j\cdot N]$ is the slice for the node $n_{j}^{\prime}$. Then we have:
$\nabla_{n_{i}}\mathcal{L}_{m}=\sum_{j=1}^{|\mathcal{N}^{\prime}|}\nabla_{n_{j}^{\prime}}\mathcal{L}_{m}\cdot W_{s_{j},s_{i}}^{post}$ (27)
Thus the second part in Equation (3) is actually the product of (i) $W_{:,s_{i}}^{post}$ the weight corresponding to the node $n_{i}$ and (ii) the derivatives of $\mathcal{L}_{m}$ w.r.t all the nodes in $H^{\prime}$, which echoes the left side of Equation (20), and the only difference is that the latter is the normalized result among all the nodes in $H$. The difference is not bothering, since activation patching compares all the contributions of nodes together, and the normalization operation exists only for convenience in the definition of the forward redundancy. Thus, the second part of Equation 3 could serve as an indicator for the forward redundancy.
Appendix F Details for the Subject-verb Disagreement Task

### v3

@@ 6 Conclusion
We describe the learning process of a model as dynamically finding the subgraph for a specific task and updating the relevant parameters in that subgraph. We propose circuit-tuning as a promising tool for the study of learning dynamics during fine-tuning. We analyze the mechanism behind learning and provide new findings and insights for our understanding of how a neural network learns and generalize. Limitations and future discussions of our work are shown in Appendix A.
7 Reproducibility Statement
To ensure reproducibility of this work, key supporting materials are distributed in the main text, appendix, and supplements. The anonymous source code for our proposed algorithm is available in Supplementary Material. Dataset details—including sources, preprocessing, and splitting—are overviewed in the main text and detailed in the Appendix and Supplementary Material. These materials enable accurate replication of our results. Finally, LLMs are used to aid our writing of this paper.
@@ Appendix A Limitations and future discussions
    The infra of our method could be further improved. We believe it is possible to apply our method to much larger models, while it requires stronger frameworks. We leave it for furture development.
  2. 2.
@@
Appendix B Transformer Architecture
@@
A single input of Transformer is $x_{0}\in\mathbb{R}^{T}$, where $T$ is the length of the sequence. The input is firstly embedded into a vector $x\in\mathbb{R}^{D\times T}$ via the embedding matrix $W_{E}\in\mathbb{R}^{D\times V}$, where $D$ is the model dimension. Then $x$ will go through $l$ layers of Transformer blocks for various processings. From the view of Elhage et al. (2021), we can think of the residual stream as a communication channel that simply receives the output of the self-attention and MLP operations. Each operation reads information from the residual stream and writes the processed information into it. Thus the residual stream is actually the linear sum of various transformations of $x$ together with the original input $x$.
@@
For the MLP operation in each layer, the input $x$ is projected into $x_{in}^{i}\in\mathbb{R}^{D_{mlp}\times T}$ via $W_{in}\in\mathbb{R}^{D_{mlp}\times D}$, and projected back to $MLP_{i}(x)\in\mathbb{R}^{D\times T}$ via $W_{out}\in\mathbb{R}^{D\times D_{mlp}}$. In the Llama architecture Touvron et al. (2023), the input $x$ is firstly projected into $x_{pre}^{i}\in\mathbb{R}^{D_{mlp}\times T}$ via $W_{gate}^{i}\in\mathbb{R}^{D_{mlp}\times D}$ and is applied with an activation layer, then a dot product is performed between the activations and $x_{in}^{i}\in\mathbb{R}^{D_{mlp}\times T}$ which is the input projected by $W_{in}\in\mathbb{R}^{D_{mlp}\times D}$. Note that we can also split the MLP into MLP heads, which is done on Llama series models in the complex tasks in our experiments. For details, please refer to Appendix E.2.
@@
Appendix C The Derivation of Edge Contribution
Similar to the definition of attribution score in Syed et al. (2023), we can define the contribution of an edge $e:n_{1}\rightarrow n_{2}$ with $n_{1}$ as the upstream node and $n_{2}$ as the downstream node. We use the indirect effect $IE$ to measure the change in the output caused by the patching of the edge. Thus given a dataset $\mathcal{X}$ for patching, the contribution of edge $e$ can be expressed as follows:
@@ Appendix C The Derivation of Edge Contribution
$=\mathbb{E}_{x_{i}\sim\mathcal{X}}\big[|IE(e;x_{i})|\big]\$(3)
Note that the contribution of edge $e$ is directly reflected in the change of the final output (the logit of the language model, etc.), which is the indirect effect caused by the change of the value in the downstream node $n_{2}$, while the change of the node $n_{2}$ is directly caused by the change of the upstream node $n_{1}$. The difference between the direct effect and the indirect effect is that the former keeps all the other nodes that could influence $n_{2}$ unchanged and only studies the influence from $n_{1}$ to $n_{2}$, while the latter allows all the changes in nodes between $n_{2}$ and the logit. For more refined definitions for these two concepts, please refer to Pearl (2001).
@@
To further simplify Equation C, we apply another Taylor expansion at $n_{1}=n_{1}(x)$ to $n_{2}^{n_{1}}$. Then we have
@@
Thus we come to the final form of edge contribution in Equation C. Our derivation takes the simplest situation into consideration, while it works well in practice. For more discussions on the relevant topic, please refer to Kramár et al. (2024). For implementation details, please refer to our code.
Appendix D Details for the Subject-verb Disagreement Task